5b+2(b+1)=16 can you help solve the equation. What is the formula?: 5b+2(b+1)=16 can you help solve the equation. What is the formula?
Answer by oscargut(682)

5b+2(b+1)=16
then
5b+2b+2=16
then
7b+2=16
then
7b=14
then
b=2
Question 154685: 5b+2(b+1)=16 can you help solve the equation. What is the formula?: 5b+2(b+1)=16 can you help solve the equation. What is the formula?
Answer by nerdybill(1283)

When you are "solving" something like:
5b+2(b+1)=16
.
What they are asking you to do is to "isolate" the unknown (in this case, b) to one side of the equation -- doesn't matter which side.
.
5b+2(b+1)=16
Start by distributing the 2 to each term inside the parentheses:
5b+(2b+2)=16
Combining like-terms on the left:
7b+2 = 16
Subtracting 2 from both sides:
7b = 14
Finally, divide both sides by 7:
b = 2 (this is your answer)